OpenStudy (anonymous):

using the substitution method, solve the system of linear equations. 3x-5y=16 4x+2y=4

OpenStudy (thomas5267):

Rearrange the second equation.

OpenStudy (welshfella):

Ok so first find y in terms of x from the second equation 2y = 4 - 4x divide through by 2 y = 2 - 2x

OpenStudy (welshfella):

now replace y is first equation by this expression for y :- 3x - 5(2 - 2x) = 16 now solve this for x